OpenStudy (anonymous):

for what value of n is n+1P3= nP4

OpenStudy (anonymous):

\[P(n+1,3)=P(n,4)\]\[\frac{(n+1)!}{(n+1-3)!}=\frac{n!}{(n-4)!}\]\[\frac{(n+1)!}{(n-2)!}=\frac{n!}{(n-4)!}\]\[(n-1).n.(n+1)=(n-3).(n-2).(n-1).n\]solve for \(n\)